find the inverse of the function y=-1/3x+2
please help me and show the work please thank you

Answer by stanbon(75887)

find the inverse of the function y=-1/3x+2
-----------------------------
Interchange x and y to get:
x = (-1/3)y + 2
---
Solve for "y":
(-1/3)y = x-2
y = -3(x-2)
y = -3x+6 (that is the inverse)
===================================
Cheers,
Stan H.